# SysCaps: Language Interfaces for Simulation Surrogates of Complex Systems

### v1

@@ 6 Experiments
(a) building-hourly
(b) stock-annual
Figure 3: Accuracy. Lower NRMSE is better. Building-hourly is the NRMSE normalized per building and per hour. Stock-annual sums over all buildings and hours before normalizing by the total consumption of the building stock. SSM/X not shown in (b) for clarity ($\sim$ 0.52 NRMSE).
Table 1: Caption quality. We estimate presence of each attribute in a SysCaps, measured by the average test accuracy of a multi-class classifier trained to predict each categorical attribute. Our metric suggests $\sim$9-12of attributes are missing or incorrect per SysCaps, due to errors made by llama-2-7b-chat.
@@
Short 88.90
Medium 90.90
Long 90.38
@@
Table 3: Regression with attribute synonyms. Pretrained language embeddings make our model robust to certain types of caption paraphrasing, such as the use of attribute synonyms. We show the difference in NRMSE between an unmodified caption and a modified caption (mean and std. across 3 random seeds), where one modified caption replaces the building type with a synonym (column 3) and the other removes the building type attribute as a baseline (column 4). A difference in means larger than the std. dev. is bolded, otherwise we use an underline.
@@
Buildings type Synonym W/ Synonym W/out Building Type FullServiceRestaurant FineDiningRestaurant 0.52$\pm$ 0.05 0.93$\pm$ 0.01 RetailStripmall ShoppingCenter 0.01$\pm$ 0.00 0.68$\pm$ 0.02 Warehouse StorageFacility 0.35$\pm$ 0.30 0.55$\pm$ 0.31 RetailStandalone ConvenienceStore 0.00$\pm$ 0.01 0.30$\pm$ 0.04 SmallOffice Co-WorkingSpace 0.03$\pm$ 0.01 0.02$\pm$ 0.02 PrimarySchool ElementarySchool 0.00$\pm$ 0.01 0.38$\pm$ 0.02 MediumOffice Workplace 0.08$\pm$ 0.02 0.03$\pm$ 0.04 SecondarySchool HighSchool -0.01$\pm$ 0.04 0.52$\pm$ 0.06 Outpatient MedicalClinic 0.02$\pm$ 0.01 0.56$\pm$ 0.09 QuickServiceRestaurant FastFoodRestaurant 0.10$\pm$ 0.07 0.83$\pm$ 0.01 LargeOffice OfficeTower 0.12$\pm$ 0.13 0.23$\pm$ 0.03 LargeHotel Five-Star Hotel 0.03$\pm$ 0.01 0.46$\pm$ 0.06 SmallHotel Motel 0.26$\pm$ 0.07 0.88$\pm$ 0.07 Hospital HealthcareFacility 0.03$\pm$ 0.04 0.62$\pm$ 0.12
Figure 4: System captions unlock text-prompt-style surrogate modeling for complex systems. We show building stock daily load profiles aggregated for Warehouse building type, created with caption templates. In tabular regression, a critical missing attribute (building square footage) can have an outsized impact on accuracy. In column two, the accuracy has significantly improved (see Figure 6 for the quantitative analysis).
Figure 5: The stock annual NRMSE decreases as missing attributes are added to a caption template. Performance improves greatly when the square footage is known.
Figure 6: Sensitivity analysis with natural language SysCap templates. a) The model has learned a physically plausible relationship between building square footage and number of stories. b) Failure case: When tested on unseen values of sqft (blue crosses), the model severely underestimates the energy consumption.
Setup: Our main experiments focus on training building stock surrogate models for the building energy simulator EnergyPlus [8]. Given an annual hourly weather timeseries ($T$ = 8,760) with 7 variables and a list of tabular building attributes, surrogates predict the building’s energy consumption at each hour of the year. Each building initially has 17 attributes; using RFE with a tuned LightGBM [28] model, we selected the 13 most important attributes. We use the commercial building split of the Buildings-900K dataset [15], which are building stock simulation runs for all commercial buildings in the United States. Since this dataset only has the energy timeseries, we extracted the building configuration and weather timeseries from the End-Use Load Profiles database [55] for each building. Our training set is comprised of 330K buildings, and we use 100 buildings for validation, and 6K held-out buildings for testing. We also reserved a held-out set of 10K buildings for RFE. We carefully tune the hyperparameters of all models (details in the Appendix).
We created three SysCaps datasets: a “medium” caption length dataset where <NUM> $\coloneqq$4-6”, a “short” dataset using 2-3 sentences and a “long” dataset using 7-9 sentences. The SSMs in our experiments are trained with medium captions. Generating these datasets with llama-2-7b-chat used $\sim$1.5K GPU hours on a cluster with 16 NVIDIA A100-40GB GPUs.
Evaluating SysCaps quality: The LLM that generates natural language SysCaps may erroneously ignore or hallucinate attributes, which can negatively impact downstream performance. We evaluate the generated captions by estimating the fraction of attributes which the LLM successfully includes per caption. To compute this metric, we train a multi-class classifier to predict each categorical attribute in a SysCaps from its text embedding. The rate of missing or incorrect attributes is around 9-12% across the “short”, “medium”, and “long” caption types, with “short” captions having the highest error. This increases our confidence that our LLM-based approach for generating natural language SysCaps preserves sufficient information for surrogate modeling.
@@ 6.1 Accuracy On Held-Out Systems
Following Emami et al. [15], we use the normalized root mean square error (NRMSE) metric to compare model accuracy, averaged across 3 random seeds. In addition to comparing the ResNet MLP, LSTM, and SSM, we also trained a tuned LightGBM baseline. Does the sequential architecture matter? Yes—Figure 3 shows that the LSTM and SSM encoders outperform both the ResNet and our carefully tuned LightGBM baseline, and the SSM outperforms the LSTM. How do different system attribute encodings compare? First, we ablate the importance of encoding the system attributes by training an SSM baseline with these inputs removed (SSM/X); this model is unable to learn this task. Surprisingly, the SSM with text templates achieves comparable test accuracy on held-out systems to the SSM with one-hot inputs. We initially expected to see a non-negligible drop in regression accuracy even for text template inputs, because the text encoder compresses the caption into a single embedding vector; however, the DistilBERT encoder is sufficiently expressive to mitigate this. The SSM with natural language SysCaps has slightly worse accuracy than the LSTM model with key-value SysCaps, yet comfortably outperforms the non-sequential models, including LightGBM. We believe the performance gap between key-value templates and natural language SysCaps is mostly explained by the caption quality (Table 2). To check whether a more powerful text encoder than DistilBERT improves accuracy, we replace it with BERT. We saw only a minor difference in the building-hourly NRMSE (Fig. 3(a)), but a large improvement from 0.069 to 0.038 in stock-annual NRMSE (Fig. 3(b)), which is comparable to the SSM with one-hot inputs.
@@ 6.2 Caption Generalization
Attribute synonyms: To quantitatively evaluate the extent to which natural language SysCaps surrogates gain a level of robustness to distribution shifts such as word order changes, synonyms, or writing style [24], we created captions for the held-out systems where the “building type” attribute is replaced by a synonym. We avoid biasing the choice of building type synonym by 5-shot prompting llama-2 to suggest the synonyms. For a control, we compare the synonym caption accuracy against accuracy on a caption with the building type attribute removed. Examples and results are shown in Table 3, where for 11/13 building type synonyms the increase in NRMSE is less than 12%, while the average increase in NRMSE for the control is 54%.
6.3 Robustness to Missing Information
Figures 46 show how a SysCaps model performs with missing attribute information in the natural language caption. We did not train the model on any captions with missing attributes. To create these figures, we progressively added more information to a caption template (Figure 4). Prediction accuracy improves as more information is given; notably, there is a large jump in accuracy once the building square footage is known, which is to be expected (Figure 6). This experiment reveals an intriguing connection between SysCaps with missing system information and traditional low-fidelity mathematical surrogate models. When attribute values are unknown, such as during the early stages of a design process, the SysCaps surrogate is naturally less accurate; similarly, a low-fidelity surrogates which uses a “coarse” mathematical approximation of a complex system sacrifices accuracy for computational efficiency.
6.4 Sensitivity Analysis From Natural Language
We visualize in Figure 6 a demonstration of using a SysCaps surrogate to conduct a sensitivity analysis on two system attributes, as might be performed for an early-stage design space exploration task. We use the caption template from Section 6.3 to create captions for each test building that enumerate all combinations of the number of stories and square footage attributes, totaling 160 configurations; the entire analysis requires simulating 960K buildings. We observe that the model has indeed learned physically plausible relationships between these two attributes. The model fails to predict the energy usage for buildings over 100K square feet—such buildings are in the “long tail” of the training data distribution.
@@ 6.5 Prompt Augmentation: Wind Farm Wake
This experiment uses the Wind Farm Wake Modeling Dataset [43], made with the FLORIS simulator, to train a surrogate to predict a wind farm’s power generation in steady-state atmospheric conditions. The difficulty of this task is in modeling losses due to wake effects, given only a coarse description of the wind farm layout. There are three numeric simulator inputs $x$ specifying atmospheric conditions, and five system attributes which include categorical variables indicating wind farm shape (there are four different layout types), number of turbines, and average turbine spacing (we do not use RFE). In this dataset, there are only 500 unique system configurations (split 3:1:1 for train, val, test), although each configuration is simulated under 500 distinct atmospheric conditions.
We explore generating multiple captions for each system configuration through prompt augmentation to increase diversity. Specifically, we replace the <STYLE> tag in the prompt with phrases encouraging different description styles, e.g., with an objective tone, with an objective tone (creative paraphrasing is acceptable), to a colleague, and to a classroom. The simulation is run assuming steady-state conditions (i.e., time-independent), so we tune hyperparameters for and train the non-sequential ResNet models. The ResNet baseline with one-hot encoded attribute inputs suffers from severe overfitting (Table 4), likely due to the small number (300) of training systems, whereas the SysCaps models generalize better to unseen systems. This suggests SysCaps can have a regularizing effect in small data settings. Notably, the prompt augmentation helps the natural language SysCaps model to achieve the lowest NRMSE.
7 Discussion
Our experiments demonstrate that, at comparable or minimal losses in accuracy with respect to standard feature engineering (e.g., one-hot encoding), real-world surrogate models can be augmented language interfaces. For a problem with only a small number of training systems available, the language inputs (with prompt augmentation) have a regularizing effect (Sec. 6.5) and actually outperformed standard feature engineering. We qualitatively and quantitatively showed that SysCaps handle missing attributes well. SysCaps unlock text-prompt-style surrogate modeling and new generalization abilities beyond what was previously possible with the standard encoding approach.
Limitations: Current BERT-style tokenizers struggle with numerical values [54]; for one example, they interpolate poorly to unseen numbers (Fig. 6). For another, because llama-2-7b-chat tends to add a comma to large numbers (e.g., $200,000$) when generating SysCaps, we found that our models failed to understand large numbers without commas (high error). Orthogonal research on improving number encodings for language model inputs [17, 57] can benefit our framework. Another potential concern is with creating SysCaps for simulators with a large (e.g., over 100) attributes. To get high quality captions, a more powerful LLM than llama-2-7b-chat may be needed.
Broader impacts: Improving the design of surrogate models for CES has the potential to accelerate the transition to cleaner energy sources. To avoid unfair outcomes from decisions made with CES surrogates, care should be taken when deciding what simulation runs to use as training data.
@@ 8 Conclusion
In this work, we introduced a learning framework for training multimodal text and timeseries surrogate models for simulations of complex energy systems such as buildings and wind farms, and described how we use LLMs to synthesize natural language descriptions of such systems, which we call SysCaps. Our findings underscore that language is a viable interface for real-world surrogate models with only minimal losses (and occasionally gains) in accuracy. Surrogates with natural language SysCaps are robust to missing system attributes and paraphrasing (e.g., synonyms of attributes).
A future extension of this work might explore how to use language to also interface with the timeseries simulator inputs, possibly through summary statistics. For example, to study how the complex system behaves when the average exogenous temperature is increased by five degrees. Moreover, an important question is how we might create surrogate foundation models that generalize not only across system configurations for a single simulator, but also generalize across different simulators.
Acknowledgments and Disclosure of Funding

### v2

@@ 6 Experiments
This section presents our experimental results, which include: quantifying the quality of LLM-generated SysCaps (Section 6.1), accuracy on held-out systems (Section 6.2), generalization under distribution shifts (Section 6.3), a design space exploration application (Section 6.4), and a study on SysCaps prompt augmentation (Section 6.5). These experiments are based on two real-world CES simulators for buildings and wind farms. All SysCaps are synthetically generated in this work. We provide additional qualitative examples of SysCaps in Appendix B.
Setup: Our main experiments focus on training building stock surrogate models for the building energy simulator EnergyPlus (Crawley et al., 2001). Given an annual hourly weather timeseries ($T$ = 8,760) with 7 variables and a list of tabular building attributes, surrogates predict the building’s energy consumption at each hour of the year. Each building initially has 17 attributes; using RFE with a tuned LightGBM (Ke et al., 2017) model, we selected the 13 most important attributes. We use the commercial building split of the Buildings-900K dataset (Emami et al., 2023b), which are building stock simulation runs representative of the United States commercial building stock. Commercial building stock surrogates can provide significant speedups compared to EnergyPlus, e.g., 96% (Zhang et al., 2021). Since this dataset only provides energy timeseries, we extract the building configuration and weather timeseries from the End-Use Load Profiles database (Wilson et al., ) for each building. Our training set is comprised of 330K buildings, and we use 100 buildings for validation, and 6K held-out buildings for testing. We also reserved a held-out set of 10K buildings for RFE. We carefully tune the hyperparameters of all models (details in Appendix A.2).
We created three SysCaps datasets: a “medium” caption length dataset where <NUM> $\coloneqq$4-6”, a “short” dataset using 2-3 sentences, and a “long” dataset using 7-9 sentences. The SSMs in our experiments are trained with medium captions. Generating these datasets with llama-2-7b-chat used $\sim$1.5K GPU hours on a cluster with 16 NVIDIA A100-40GB GPUs.
6.1 Evaluating caption quality
The LLM that generates natural language SysCaps may erroneously ignore or hallucinate attributes, which may negatively impact downstream performance. To test this, we propose evaluating generated captions by estimating the fraction of attributes which the LLM successfully includes per caption. To compute this metric, we train a multi-class classifier to predict each categorical attribute in a SysCaps from its text embedding. We used a held-out validation set of captions to check that the classifier was not overfitting. The test rate of missing or incorrect attributes is around 9-12% across the “short”, “medium”, and “long” caption types, with “short” captions having the highest error (Table 3). This increases our confidence that our LLM-based approach for generating natural language SysCaps preserves sufficient information for surrogate modeling.
Table 1: Accuracy. We show the mean NRMSE across 3 random seeds. Lower NRMSE is better. Building-hourly is the NRMSE normalized per building and per hour. Stock-annual first sums the predictions and targets over all buildings and hours, before computing the NRMSE (equivalent to the normalized mean bias error—see Appendix A.1). When using a BERT encoder instead of DistilBERT, SysCaps surrogates achieve better accuracy than one-hot baselines. We trained one SSM without any attribute information (Attribute Encoding “X”) as an ablation study; the poor accuracy shows that our multimodal architecture successfully learns to fuse the text-based attribute and timeseries inputs.
Model Text Encoder Attribute Encoding Buildings-Hourly (NRMSE) Stock-Annual (NRMSE) SSM - X 1.712$\pm$ 0.003 0.658$\pm$ 0.008 SSM - onehot 0.450$\pm$ 0.019 0.041$\pm$ 0.021 LSTM - onehot 0.449$\pm$ 0.025 0.045$\pm$ 0.024 ResNet - onehot 0.634$\pm$ 0.009 0.072$\pm$ 0.008 LightGBM - onehot 0.679$\pm$ 0.014 0.094$\pm$ 0.003 SSM DistilBERT SysCaps-nl 0.532$\pm$ 0.010 0.069$\pm$ 0.010 SSM BERT SysCaps-nl 0.543$\pm$ 0.011 0.035$\pm$ 0.005 SSM DistilBERT SysCaps-kv 0.454$\pm$ 0.012 0.046$\pm$ 0.003 SSM BERT SysCaps-kv 0.450$\pm$ 0.007 0.020$\pm$ 0.012 LSTM DistilBERT SysCaps-kv 0.489$\pm$ 0.021 0.063$\pm$ 0.005 LSTM BERT SysCaps-kv 0.439$\pm$ 0.037 0.022$\pm$ 0.011 ResNet DistilBERT SysCaps-kv 0.633$\pm$ 0.020 0.081$\pm$ 0.011 ResNet BERT SysCaps-kv 0.670$\pm$ 0.043 0.049$\pm$ 0.015
Figure 3: System captions unlock text-prompt-style surrogate modeling for complex systems. We show building stock daily load profiles aggregated for Warehouse building type, created with caption templates. From left to right, we use captions with one, three, and six attributes.
Table 2: Caption quality. We estimate the presence of each attribute in a SysCap, measured by the average test accuracy of a multi-class classifier trained to predict each categorical attribute. Our metric suggests $\sim$9-12of attributes are missing or incorrect per SysCap, due to errors made by llama-2-7b-chat.
@@ 6.1 Evaluating caption quality
Short (2-3 sent.) 88.90
Medium (4-6 sent.) 90.90
Long (7-9 sent.) 90.38
@@
Table 4: Generalization to attribute synonyms. We quantify how text embeddings make our models robust to the use of attribute synonyms. The metric is the difference in NRMSE between the original caption and the modified caption. We bold the best method, i.e, closest to 0. Column 3 replaces the building type with a synonym, column 4 removes the building type and sub-type attributes from the caption, and column 5 randomly swaps the correct building type attributes with incorrect ones.
@@
Building Type Synonym With Synonym Without Building Type Random Building Type FullServiceRestaurant FineDiningRestaurant 0.52$\pm$ 0.05 0.93$\pm$ 0.01 1.17$\pm$ 0.07 RetailStripmall ShoppingCenter 0.01$\pm$ 0.00 0.68$\pm$ 0.02 0.28$\pm$ 0.04 Warehouse StorageFacility 0.35$\pm$ 0.30 0.55$\pm$ 0.31 4.02$\pm$ 0.32 RetailStandalone ConvenienceStore 0.00$\pm$ 0.01 0.30$\pm$ 0.04 0.40$\pm$ 0.03 SmallOffice Co-WorkingSpace 0.03$\pm$ 0.01 0.02$\pm$ 0.02 1.95$\pm$ 0.30 PrimarySchool ElementarySchool 0.00$\pm$ 0.01 0.38$\pm$ 0.02 0.52$\pm$ 0.17 MediumOffice Workplace 0.08$\pm$ 0.02 0.03$\pm$ 0.04 0.91$\pm$ 0.11 SecondarySchool HighSchool -0.01$\pm$ 0.04 0.52$\pm$ 0.06 0.67$\pm$ 0.33 Outpatient MedicalClinic 0.02$\pm$ 0.01 0.55$\pm$ 0.09 0.32$\pm$ 0.06 QuickServiceRestaurant FastFoodRestaurant 0.10$\pm$ 0.07 0.83$\pm$ 0.01 0.85$\pm$ 0.01 LargeOffice OfficeTower 0.12$\pm$ 0.13 0.23$\pm$ 0.03 0.29$\pm$ 0.11 LargeHotel Five-Star Hotel 0.03$\pm$ 0.01 0.46$\pm$ 0.06 0.29$\pm$ 0.09 SmallHotel Motel 0.26$\pm$ 0.07 0.88$\pm$ 0.07 0.70$\pm$ 0.14 Hospital HealthcareFacility 0.03$\pm$ 0.04 0.62$\pm$ 0.12 0.20$\pm$ 0.07
@@ 6.2 Accuracy On Held-Out Systems
Following Emami et al. (2023b), we use the normalized root mean square error (NRMSE) metric to compare model accuracy, averaged across 3 random seeds. In addition to comparing the ResNet MLP, LSTM, and SSM, we also trained a tuned LightGBM baseline. Does the sequential architecture matter? Yes—Table 1 shows that the LSTM and SSM encoders outperform both the ResNet and our carefully tuned LightGBM baseline, and the SSM outperforms the LSTM. How do different system attribute encoding approaches compare? Surprisingly, the best SSM and LSTM with key-value SysCaps achieve comparable building-hourly test accuracy to one-hot baselines and the best accuracy at the stock-annual aggregation level (equivalent to the normalized mean bias error—see Appendix A.1). The SSM and LSTM with natural language SysCaps has slightly worse accuracy than with the key-value SysCaps, yet they comfortably outperform the non-sequential models (including LightGBM) and the one-hot baselines at the stock-annual aggregation level. We initially expected to see a non-negligible drop in regression accuracy for SysCaps models, even key-value SysCaps, because the text encoder has to compress the caption into a single embedding vector. However, the BERT encoder is sufficiently expressive to mitigate this. We observe that the stock-level NRMSE reduces by about half when switching from DistilBERT to BERT. We believe the performance gap between key-value templates and natural language SysCaps is mostly explained by the caption quality (Table 3). We ablate the importance of the system attributes by training an SSM baseline with these inputs removed (SSM with Attribute Encoding “X”); this model is unable to learn this task.
Figure 3 qualitatively shows how a SysCaps model performs with natural language captions (created with a sentence template) provided to the model. Note that we did not train our models on any captions with missing attributes. Prediction accuracy improves as more information is given; notably, there is a large jump in accuracy once the building square footage is known.
@@ 6.3 Caption Generalization
Attribute synonyms: To quantitatively evaluate the extent to which natural language SysCaps surrogates gain a level of robustness to distribution shifts such as word order changes, synonyms, or writing style (Hendrycks et al., 2020), we created captions for the held-out systems where the “building type” attribute is replaced by a synonym. We avoid biasing the choice of building type synonym by 5-shot prompting llama-2 to suggest the synonyms. There are two baselines we compare the synonym caption accuracy against: 1) accuracy when testing the model on captions with the building type attribute removed, and 2) accuracy when testing the model on captions with a random building type. Examples and results are shown in Table 4, where for 11/14 building type synonyms the increase in NRMSE is less than 13%, while the average increase for the two baselines is 54% and 90%, respectively.
6.4 Design Space Exploration Application Using Language
Figure 4: Design space exploration using language. a) We show that the model has learned a physically plausible relationship between building square footage (sqft) and number (#) of stories. b) Failure case: When tested on unseen values of sqft (blue crosses), the model’s predictions appear to be physically implausible—the model underestimates the energy consumption at these sqft values.
We visualize in Figure 4(a) a use of SysCaps to conduct a sensitivity analysis on two system attributes, as might be performed for an early-stage design space exploration task. We use a simple template to create a caption for each test building that enumerates all combinations of the number of stories and square footage attributes, totaling 160 configurations; the entire analysis requires simulating 960K buildings, and took 1 hour on a single NVIDIA A100 GPU. We observe that the model has indeed learned physically plausible relationships between these two attributes. However, the model fails to predict the energy usage for buildings over 100K square feet—such buildings are in the “long tail” of the training data distribution. Figure 4(b) also shows a failure case where the model underestimates energy usage at unseen numeric building square footage attribute values.
@@ 6.5 Prompt Augmentation: Wind Farm Wake
This experiment uses the Wind Farm Wake Modeling Dataset (Ramos et al., 2023), made with the FLORIS simulator, to train a surrogate to predict a wind farm’s power generation in steady-state atmospheric conditions. The speed-up provided by surrogate models for downstream optimization use-cases is $\sim$700$\times$ (Harrison-Atlas et al., 2024). The difficulty of this task is in modeling losses due to wake effects, given only a coarse description of the wind farm layout. There are three numeric simulator inputs $x$ specifying atmospheric conditions, and five system attributes which include categorical variables indicating wind farm shape (four different layout types), number of turbines, and average turbine spacing. We do not use RFE. In this dataset, there are only 500 unique system configurations (split 3:1:1 for train, val, test), although each configuration is simulated under 500 distinct atmospheric conditions.
@@
We explore generating multiple captions for each system configuration through prompt augmentation to increase diversity. Specifically, we replace the <STYLE> tag in the prompt with phrases encouraging different description styles, e.g., with an objective tone, with an objective tone (creative paraphrasing is acceptable), to a colleague, and to a classroom. The simulation is run assuming steady-state conditions (i.e., time-independent), so we tune hyperparameters for and train the non-sequential ResNet models. The ResNet baseline with one-hot encoded attributes suffers from severe overfitting (Table 5), likely due to the small number (300) of training systems, whereas the SysCaps models generalize better to unseen systems. This suggests SysCaps can have a regularizing effect in small data settings. Notably, the prompt augmentation helps the natural language SysCaps model to achieve the lowest NRMSE.
@@ 7 Conclusion
In this work, we introduced a lightweight, multimodal text and timeseries surrogate models for complex energy systems such as buildings and wind farms, and described a process for using LLMs to synthesize natural language descriptions of such systems, which we call SysCaps. Our experiments showcase SysCaps-augmented surrogates that achieve better accuracy than standard feature engineering (e.g., one-hot encoding) while also enjoying the advantages of using text embeddings such as robustness to caption paraphrasing (e.g., synonyms of attributes). For a problem with only a small number of training systems available, we showed that SysCaps prompt augmentation has a regularizing effect that helps mitigate overfitting. Overall, these results underscore that language is a viable interface for interacting with real-world surrogate models.
Limitations: Current BERT-style tokenizers struggle with numerical values (Wallace et al., 2019); for one example, they interpolate poorly to unseen numbers (Figure 4(b)). For another, because llama-2-7b-chat tends to add a comma to large numbers (e.g., $200,000$) when generating SysCaps, we found that our surrogates failed to understand large numbers without commas (had high error). Orthogonal research on improving number encodings for language model inputs (Golkar et al., 2023; Yan et al., 2024) can benefit our framework. Another potential concern is with creating SysCaps for simulators with a large number of attributes (e.g., over 100). A more powerful LLM than llama-2-7b-chat with a longer context window may be needed in this case. In general, we expect that more powerful LLMs will further improve the quality of the training captions.
Broader impacts: Improving surrogate models for CES has the potential to accelerate the transition to cleaner energy sources. To avoid unfair outcomes from decisions made with CES surrogates, care should be taken when deciding what simulation runs to use as training data and when selecting an LLM to use for caption generation.
Future work: A future extension of this work might explore how to use language to also interface with the timeseries simulator inputs, possibly through summary statistics. For example, to study how the complex system behaves when the average exogenous temperature is increased by five degrees. Another promising extension is to conduct a series of human studies with domain scientists to evaluate the quality of SysCaps in the context of, e.g., system design optimization. Moreover, an important question is how we might create surrogate foundation models that generalize not only across system configurations for a single simulator, but also generalize across different simulators.
Acknowledgments